Does the line below pass through the point (2,-2)?
4x + y = 10

Answer by funmath(2873)

Does the line below pass through the point (2,-2)?
4x + y = 10
:
(x,y)=(2,-2), substitute 2 in for x and -2 in for y, if the equation's sides are equal, the point is on the line. If not it's not.
:
4(2)+(-2)=10 ?
8-2=10
6 not=10
The point (2,-2) is not on the line.
Happy Calculating!!!